OpenStudy (madgirlwithabluebox):

please help? Will give medal to the person who explains his to me. Use the quadratic formula to solve the equation. –x2 + 7x = 8

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[\displaystyle\boxed{x_{1,2}=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}}\] Start by rearranging the equation into the form: \[ax^2+bx+c=0\] Then find the coefficients \(a\),\(b\),\(c\).

OpenStudy (madgirlwithabluebox):

so right now i have -1x^2+7x+?=8

OpenStudy (unklerhaukus):

that's a good start, but you need to get `= 0` on the right hand side, so take away 8 from both sides of the equation...

OpenStudy (madgirlwithabluebox):

-1x^2+7x-8=0 is this right?

OpenStudy (unklerhaukus):

Yes, so what are the coefficients?

OpenStudy (madgirlwithabluebox):

-1 and 7?

OpenStudy (unklerhaukus):

yeah a = -1 and b =7 but what is c =

OpenStudy (madgirlwithabluebox):

-8 but i thought a coefficient needed a variable so i didnt include it

OpenStudy (unklerhaukus):

oh, um, yeah ok, i guess c isn't really a coefficient, but, yes c =-8

OpenStudy (unklerhaukus):

Now just sub these; \[a =-1, \qquad b=7, \qquad c=-8\]into the Quadratic formula \[\displaystyle\boxed{x_{1,2}=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}}\]

OpenStudy (unklerhaukus):

if you want to take it slow, you could work out \[\sqrt{b^2-4ac}\] first

OpenStudy (madgirlwithabluebox):

ok i did that and have 17 under the square root symbol

OpenStudy (unklerhaukus):

very good, notice that 17 is prime, so the square root of it will not simply, i.e your final answers will have \(\sqrt{17}\) terms in them

OpenStudy (unklerhaukus):

now sub in \(a\) and \(b\) into \[x_{1,2}=\dfrac{-b\pm\sqrt{17}}{2a}\]

OpenStudy (madgirlwithabluebox):